Which of the following would require the
greatest number of calories?
1. Heating 10 g of water from 10◦C to 40◦C
2. Heating 1000 g of water from 10◦C to 12◦C
3. Heating 100 g of water from 10◦C to 50◦C
4. Heating 1 g of water from 10◦C to 80◦C
1
Expert's answer
2020-04-02T10:39:14-0400

1)


"Q=(10)(40-10)=300\ calories"

2)


"Q=(1000)(12-10)=2000\ calories"

3)


"Q=(100)(50-10)=4000\ calories"

4)


"Q=(1)(80-10)=70\ calories"

Answer: 3. Heating 100 g of water from 10◦C to 50◦C